# Michael Witbrock1 Jiamou Liu1 1

### v4

@@
Qiming Bao1, 2, Alex Yuxuan Peng1, Zhenyun Deng3, Wanjun Zhong4, Gaël Gendron1,
Timothy Pistotti1, Neşet Tan1, Nathan Young1, Yang Chen1, Yonghua Zhu1, Paul Denny5,
Michael Witbrock1
@@
3Department of Computer Science, University of Cambridge, The United Kingdom
@@ Abstract
Combining large language models with logical reasoning enhances their capacity to address problems in a robust and reliable manner. Nevertheless, the intricate nature of logical reasoning poses challenges to gathering reliable data from the web for building comprehensive training datasets, subsequently affecting the performance on downstream tasks. To address this, we introduce a novel logic-driven data augmentation approach, AMR-LDA. AMR-LDA converts the original text into an Abstract Meaning Representation (AMR) graph, a structured semantic representation that encapsulates the logic structure of the sentence, upon which operations are performed to generate logically modified AMR graphs. The modified AMR graphs are subsequently converted back into text to create augmented data. Notably, our methodology is architecture-agnostic and enhances both generative large language models, such as GPT-3.5 and GPT-4, through prompt augmentation, and discriminative large language models through contrastive learning with logic-driven data augmentation. Empirical evidence underscores the efficacy of our proposed method with improvement in performance across seven downstream tasks, such as reading comprehension requiring logical reasoning, textual entailment, and natural language inference. Furthermore, our method leads on the ReClor leaderboard [^1] https://eval.ai/web/challenges/challenge-page/503/leaderboard/1347. The source code and data are publicly available [^2] https://bit.ly/3OWKe8r.
@@
Qiming Bao1, 2, Alex Yuxuan Peng1, Zhenyun Deng3, Wanjun Zhong4, Gaël Gendron1, Timothy Pistotti1, Neşet Tan1, Nathan Young1, Yang Chen1, Yonghua Zhu1, Paul Denny5, Michael Witbrock1,  and Jiamou Liu1 1Strong AI Lab, NAOInstitute, Waipapa Taumata Rau - The University of Auckland 2Xtracta, New Zealand 3Department of Computer Science, University of Cambridge, The United Kingdom 4School of Computer Science and Engineering, Sun Yat-Sen University, China 5School of Computer Science, The University of Auckland, New Zealand {qbao775,ypen260,ntan607,yche767,ggen187}@aucklanduni.ac.nz, zd302@cam.au.uk
@@ 1 Introduction
Enabling pre-trained large language models (LLMs) to reliably perform logical reasoning is an important step towards strong artificial intelligence Chollet (2019). However, data annotation for logical reasoning tasks is a difficult, time-consuming and costly process that has led to the scarcity of large-scale logical reasoning datasets derived from natural language on the web. Therefore, LLMs are usually trained on generic corpora or smaller datasets on logical reasoning that lead to poor generalisation Wang et al. (2022). The automatic augmentation of logical reasoning data has the potential to enhance the generalisation and performance of LLMs on logical reasoning tasks.
To address this challenge, we propose a logic-driven data augmentation method based on Abstract Meaning Representation (AMR). AMR is a structural representation of the semantic meaning and logical structure of a text via a rooted directed acyclic graph (DAG) Shou et al. (2022). Figure 1 shows an example of AMR graph. The AMR graph can be easily modified by changing the nodes or arguments to create logically equivalent or nonequivalent graphs. By taking advantage of the ease of logical manipulation of AMR graph and the end-to-end conversion between natural language and AMR graph, our proposed data augmentation is not task-specific or template-dependent, and can generate logically equivalent and nonequivalent sentences that are diverse in language.
@@
In order to improve the performance of LLMs on downstream tasks that require logical reasoning, we investigate two different applications of the proposed logic-driven data augmentation for two different types of language models. In this paper, we call models such as RoBERTa Liu et al. (2019) and DeBERTa He et al. (2021) as discriminative large language models, and models like GPT-3.5 OpenAI (2023a) as generative LLMs. We improve the reasoning ability of discriminative large language models by applying contrastive learning to identify logically equivalent and nonequivalent sentence pairs generated using the proposed data augmentation, before fine-tuning the model further on downstream tasks. In order to improve the performance of generative LLMs on logical reasoning tasks without fine-tuning, we augment the input prompt by extending the question context and options using data augmentation. We summarize key contributions as follows:
@@ 2 Related Work
Logical reasoning is rigorous thinking to derive a conclusion based on a given premise Seel (2011); Bronkhorst et al. (2020). Existing datasets on reasoning can be categorised into two levels: sentence level, including tasks like natural language inference that assess if one sentence logically follows from another (e.g., MNLI Williams et al. (2018), RTE Wang et al. (2018), MRPC Dolan and Brockett (2005), QNLI Rajpurkar et al. (2016), QQP Wang et al. (2018)); passage level, which requires logical deduction from given contexts, questions, and multiple choices (e.g., PARARULE Clark et al. (2021), PARARULE-Plus Bao et al. (2022)) and reading comprehension tasks (e.g., ReClor Yu et al. (2020), LogiQA Liu et al. (2021)). We introduce an abstract meaning representation-based methodology for logic-driven data augmentation aimed at enhancing models’ logical reasoning capabilities across these tasks.
@@ 3.2 AMR-Based Logic-Driven Data Augmentation
Alan is kind. $\Leftrightarrow$ Alan is not unkind.
We apply the double negation law only to those sentences and their AMR graphs that do not contain the “:polarity -” argument which represents negative polarity. There are several words that can be represented as “:polarity -”, such as “not”, “no”, “never”, “none”, and “nothing”. A representative example can be seen in Table 1 and in Appendix Figure 8. The original sentence is “The bald eagle is strong”. The logically equivalent sentence we construct using double negation law is “The bald eagle is not weak”, while the logically nonequivalent sentence is “The bald eagle is weak”. Note that the generated sentences do not contain the word “not” twice. We avoid generating sentences with “not” appearing multiple times consecutively because they are uncommon and unnatural. The process of applying double negation law is as follows: convert the sentence into an AMR graph; augment the AMR graph by adding a negative polarity argument “: polarity -”; convert the modified AMR graph back into a natural language sentence; lastly, replace the adjective word with its antonym by using WordNet Miller (1992). To create logically nonequivalent sentences, we randomly delete or add a negative polarity argument “:polarity -” in the AMR graph. Additionally, we randomly sample a sentence from the corpus and consider it as logically nonequivalent to the original sentence.
@@ 4.1 Datasets
ReClor Yu et al. (2020) and LogiQA Liu et al. (2021) are two challenging logical reasoning datasets. ReClor is collected from Graduate Management Admission Test (GMAT) and Law School Admission Test (LSAT). LogiQA is collected by National Civil Service Examination Liu et al. (2021). Additionally, we performed evaluation on five datasets for natural language inference and textual entailment tasks: MNLI Williams et al. (2018), RTE Wang et al. (2018), MRPC Dolan and Brockett (2005), QNLI Rajpurkar et al. (2016), and QQP Wang et al. (2018). MNLI, RTE, and MRPC assess the relationship between two sentences, while QNLI focuses on the relationship between a question and a sentence, and QQP evaluates the relationship between two questions.
@@
In this paper, we perform contrastive learning for discriminative large language models on sentences augmented from a synthetic dataset. This dataset contains 14,962 sentences with different combinations of 23 entities, 2 relations and 40 attributes. Synthetic data is used to generate more controllable logical sentences. More details about the synthetic dataset can be found in the Appendix Section E.
@@ 4.2 Settings
We conduct all the experiments on 8 NVIDIA A100 GPUs, each with 80G of VRAM. We conducted the primary experiments on the ReClor and LogiQA datasets using three different random seeds, and reported the average values in Table 2. We adopted the parse_xfm_bart_large and T5Wtense models from AMRLib [^3] https://amrlib.readthedocs.io/en/latest/models/ to perform text-to-AMR and AMR-to-text conversions when generating logically augmented sentence pairs. The reason why we select those two models is explained in subsection C. In our experiments, we use RoBERTa Liu et al. (2019) and DeBERTa He et al. (2021) as the discriminative large language models. We also compare our method against MERIt Jiao et al. (2022) and IDoL Xu et al. (2023), the leading models on the ReClor leaderboard. As for generative large language models, we utilise GPT-3.5 (gpt-3.5-turbo) OpenAI (2023a) and GPT-4 OpenAI (2023b). More details about the experiment, case studies and confidence intervals can be found in Appendix Section BDD.1, and F.
@@ 4.4 Prompt Augmentation for Generative LLM
GPT-3.5 AMR-LDA performs better than GPT-3.5 on the general test set, which includes both test-E and test-H. The ReClor test set is hidden, so we do not have access to the detailed results for test-E and test-H. Therefore, we cannot provide a clear explanation as to why AMR-LDA seems to decrease the test-H metric for GPT-3.5. However, a detailed examination of the results reveals that GPT-3.5 achieves only a 0.5375 test accuracy on test-H, whereas GPT-4 attains a 0.8857 test accuracy on the same test. Furthermore, GPT-4 with AMR-LDA performs better on all the ReClor and LogiQA test sets. This suggests that GPT-3.5 might not be as effective in comprehending complex logical reasoning as GPT-4 and GPT-3.5 may struggle with understanding augmented prompts.
@@ 5 Conclusion
The sparsity of web data related to logical reasoning constrains the advancement of large language models in their performance on logical reasoning tasks. Existing methods for constructing logically equivalent sentences are restricted to templates and specific datasets. Our AMR-LDA considers more logical equivalence laws than existing methods do, and it does not reply on any ad-hoc templates. We apply AMR-LDA to fine-tuning discriminative LLMs and prompt augmentation of generative LLMs (GPT-3.5 and GPT-4), which yield better results than baseline methods on logical reasoning tasks.
@@ 6 Human Evaluation
Human evaluation is conducted to evaluate the correctness and fluency of the logically manipulated sentences generated using AMR-LDA and LReasoner-LDA. We constructed the survey with 20 questions, and each question consists of two randomly selected sentences: one from those generated by our AMR-LDA and the other by LReasoner-LDA, respectively. 45 participants participate to the survey anonymously. We ask them to evaluate the sentences on two aspects: which sentence is logically equivalent to the original sentence or both of them are logically equivalent to the original sentence. The other one is which sentence is more fluent. 63.92% and 76.44% of people prefer AMR-LDA’s logically equivalent and fluent sentences over those generated by LReasoner-LDA.
@@ 8 Ethics Statement
All the data used in this paper are either synthetically generated or open-source datasets. All the code used to run the experiments is written using open-source libraries or adapted from published code from other papers. We will also release our code and any synthetically generated data to ensure that the work can be reproduced. The human evaluation has been approved by the Ethics Committee of the main authors’ employer.
@@ Appendix D Case Studies
      AMR-LDA       Lions aren’t funny
      or Tigers beautiful.

### v6

@@
Qiming Bao, Alex Yuxuan Peng, Zhenyun Deng, Wanjun Zhong, Gael Gendron,
Timothy Pistotti, Neset Tan, Yang Chen, Yonghua Zhu,
Paul Denny, Michael Witbrock,
   Jiamou Liu
Address line
Address line
   Qiming Bao1, 2, Alex Yuxuan Peng1, Zhenyun Deng3, Wanjun Zhong4,
Gaël Gendron1, Timothy Pistotti1, Neşet Tan1, Nathan Young1, Yang Chen1,
Yonghua Zhu1, Paul Denny5, Michael Witbrock1
@@
3Department of Computer Science and Technology, University of Cambridge, UK
@@ Abstract
Combining large language models with logical reasoning enhances their capacity to address problems in a robust and reliable manner. Nevertheless, the intricate nature of logical reasoning poses challenges when gathering reliable data from the web to build comprehensive training datasets, subsequently affecting performance on downstream tasks. To address this, we introduce a novel logic-driven data augmentation approach, AMR-LDA. AMR-LDA converts the original text into an Abstract Meaning Representation (AMR) graph, a structured semantic representation that encapsulates the logical structure of the sentence, upon which operations are performed to generate logically modified AMR graphs. The modified AMR graphs are subsequently converted back into text to create augmented data. Notably, our methodology is architecture-agnostic and enhances both generative large language models, such as GPT-3.5 and GPT-4, through prompt augmentation, and discriminative large language models through contrastive learning with logic-driven data augmentation. Empirical evidence underscores the efficacy of our proposed method with improvement in performance across seven downstream tasks, such as reading comprehension requiring logical reasoning, textual entailment, and natural language inference. Furthermore, our method leads on the ReClor leaderboard [^1] https://eval.ai/web/challenges/challenge-page/503/leaderboard/1347. The source code and data are publicly available [^2] AMR-LDA GitHub Repository.
@@
Qiming Bao1, 2, Alex Yuxuan Peng1, Zhenyun Deng3, Wanjun Zhong4, Gaël Gendron1, Timothy Pistotti1, Neşet Tan1, Nathan Young1, Yang Chen1, Yonghua Zhu1, Paul Denny5, Michael Witbrock1,  and Jiamou Liu1 1Strong AI Lab, NAOInstitute, Waipapa Taumata Rau - The University of Auckland 2Xtracta, New Zealand 3Department of Computer Science and Technology, University of Cambridge, UK 4School of Computer Science and Engineering, Sun Yat-Sen University, China 5School of Computer Science, The University of Auckland, New Zealand {qbao775,ypen260,ntan607,yche767,ggen187}@aucklanduni.ac.nz, zd302@cam.au.uk
@@ 1 Introduction
Enabling pre-trained large language models (LLMs) to reliably perform logical reasoning is an important step towards strong artificial intelligence Chollet (2019). However, data annotation for logical reasoning tasks is a difficult, time-consuming and costly process that has led to the scarcity of large-scale logical reasoning datasets derived from natural language on the web. Therefore, LLMs are usually trained on generic corpora or smaller logical reasoning datasets that lead to poor generalisation Wang et al. (2022). Automatic augmentation of logical reasoning data has the potential to enhance the generalisation and performance of LLMs on logical reasoning tasks.
To address this challenge, we propose a logic-driven data augmentation method based on Abstract Meaning Representation (AMR). AMR is a structural representation of the semantics and logical structure of text via a rooted directed acyclic graph (DAG) Shou et al. (2022). Figure 1 shows an example of an AMR graph. The AMR graph can be easily modified by changing nodes or arguments to create logically equivalent or nonequivalent graphs. By taking advantage of the ease of logical manipulation of AMR graph and of end-to-end conversion between natural language and AMR graphs, our proposed data augmentation is not task-specific or template-dependent, and can generate logically equivalent and nonequivalent sentences that are diverse in their use of language.
@@
In order to improve the performance of LLMs on downstream tasks requiring logical reasoning, we investigate two different applications of the proposed logic-driven data augmentation for two different types of language models. In this paper, we describe models such as RoBERTa Liu et al. (2019) and DeBERTa He et al. (2021) as discriminative large language models, and models like GPT-3.5 OpenAI (2023a) as generative LLMs. We improve the reasoning ability of discriminative large language models by applying contrastive learning to identify logically equivalent and nonequivalent sentence pairs generated using the proposed data augmentation before fine-tuning the model further on downstream tasks. In order to improve the performance of generative LLMs on logical reasoning tasks without fine-tuning, we augment the input prompt by extending the question context and options using data augmentation. We summarize the paper’s key contributions as follows:
@@ 2 Related Work
Logical reasoning is rigorous thinking to derive a conclusion based on a given premise Seel (2011); Bronkhorst et al. (2020). Existing reasoning datasets reasoning can be categorised into two levels: sentence level, including tasks like natural language inference that assess if one sentence logically follows from another (e.g., MNLI Williams et al. (2018), RTE Wang et al. (2018), MRPC Dolan and Brockett (2005), QNLI Rajpurkar et al. (2016), QQP Wang et al. (2018)); passage level, which requires logical deduction from given contexts, questions, and multiple choices (e.g., PARARULE Clark et al. (2021), PARARULE-Plus Bao et al. (2022)) and reading comprehension tasks (e.g., ReClor Yu et al. (2020), LogiQA Liu et al. (2021)). We introduce an abstract meaning representation-based methodology for logic-driven data augmentation aimed at enhancing models’ logical reasoning capabilities across these tasks.
@@ 3.2 AMR-Based Logic-Driven Data Augmentation
It is raining. $\Leftrightarrow$ It is not the case that it is not raining.
We apply the double negation law only to those sentences and their AMR graphs that do not contain the “:polarity -” argument which represents negative polarity. There are several words that can be represented as “:polarity -”, such as “not”, “no”, “never”, “none”, and “nothing”. A representative example can be seen in Table 1 and in Appendix Figure 8. The original sentence is “The bald eagle is strong”. The logically equivalent sentence we construct using the double negation law is “The bald eagle is not weak”, while the logically nonequivalent sentence is “The bald eagle is weak”. Note that the generated sentences do not contain the word “not” twice. We avoid generating sentences with “not” appearing multiple times consecutively because they are uncommon and unnatural. The process of applying double negation law is as follows: convert the sentence into an AMR graph; augment the AMR graph by adding a negative polarity argument “: polarity -”; convert the modified AMR graph back into a natural language sentence; lastly, replace the adjective word with its antonym by using WordNet Miller (1992). To create logically nonequivalent sentences, we randomly delete or add a negative polarity argument “:polarity -” in the AMR graph. Additionally, we randomly sample a sentence from the corpus and consider it as logically nonequivalent to the original sentence.
@@ 4.1 Datasets
ReClor Yu et al. (2020) and LogiQA Liu et al. (2021) are two challenging logical reasoning datasets. ReClor is collected from the Graduate Management Admission Test (GMAT) and Law School Admission Test (LSAT). LogiQA is collected from the National Civil Service Examination Liu et al. (2021). Additionally, we performed evaluation on five datasets for natural language inference and textual entailment tasks: MNLI Williams et al. (2018), RTE Wang et al. (2018), MRPC Dolan and Brockett (2005), QNLI Rajpurkar et al. (2016), and QQP Wang et al. (2018). MNLI, RTE, and MRPC assess the relationship between two sentences, while QNLI focuses on the relationship between a question and a sentence, and QQP evaluates the relationship between two questions.
@@
In this paper, we performed contrastive learning for discriminative large language models on sentences augmented from a synthetic dataset. This dataset contains 14,962 sentences with different combinations of 23 entities, 2 relations and 40 attributes. Synthetic data was used to generate more controllable logical sentences. More details about the synthetic dataset can be found in the Appendix Section E.
@@ 4.2 Settings
All experiments were conducted on 8 NVIDIA A100 GPUs, each with 80G of VRAM. Primary experiments on the ReClor and LogiQA datasets used three different random seeds; the average values are reported in Table 2. The parse_xfm_bart_large and T5Wtense models from AMRLib [^3] https://amrlib.readthedocs.io/en/latest/models/ were used for text-to-AMR and AMR-to-text conversions when generating logically augmented sentence pairs. The reason for selecting those two models is explained in subsection C. In our experiments,RoBERTa Liu et al. (2019) and DeBERTa He et al. (2021) were used as the discriminative large language models. We also compared our method against MERIt Jiao et al. (2022) and IDoL Xu et al. (2023), the leading models on the ReClor leaderboard. As for generative large language models, we applied GPT-3.5 (gpt-3.5-turbo) OpenAI (2023a) and GPT-4 OpenAI (2023b). More details about the experiments, case studies and confidence intervals can be found in Appendix Section BDD.1, and F.
@@ 4.4 Prompt Augmentation for Generative LLM
GPT-3.5 AMR-LDA performs better than GPT-3.5 on the general test set, which includes both test-E and test-H. The ReClor test set is hidden, so we do not have access to the detailed results for test-E and test-H. Therefore, we cannot provide a clear explanation as to why AMR-LDA seems to decrease the test-H metric for GPT-3.5. However, a detailed examination of the results reveals that GPT-3.5 achieves only a 0.5375 test accuracy on test-H, whereas GPT-4 attains a 0.8857 test accuracy on the same test. Furthermore, GPT-4 with AMR-LDA performs better on all the ReClor and LogiQA test sets. This suggests that GPT-3.5 might not be as effective in comprehending complex logical reasoning as GPT-4 and GPT-3.5 may understand augmented prompts poorly.
@@ 5 Conclusion
The sparsity of web data related to logical reasoning constrains the advancement of large language models in their performance on logical reasoning tasks. Existing methods for constructing logically equivalent sentences had been restricted to templates and specific datasets. Our AMR-LDA considers more logical equivalence laws than existing methods do, and it does not reply on any ad-hoc templates. We applied AMR-LDA to fine-tuning discriminative LLMs and prompt augmentation of generative LLMs (GPT-3.5 and GPT-4), yielding better results than baseline methods on logical reasoning tasks.
@@ 6 Human Evaluation
Human evaluation was conducted to evaluate the correctness and fluency of the logically manipulated sentences generated using AMR-LDA and LReasoner-LDA. We constructed a survey with 20 questions, each question consisting of two randomly selected sentences: one from those generated by our AMR-LDA and the other by LReasoner-LDA. 45 participants completed the survey anonymously. We asked them to evaluate the sentences in two aspects: 1) which sentence is logically equivalent to the original sentence, or whether both of them are logically equivalent to the original sentence, and 2) which sentence is more fluent. 63.92% and 76.44% of people prefered AMR-LDA’s logically equivalent and fluent sentences over those generated by LReasoner-LDA.
@@ 8 Ethics Statement
All the data used in this paper are either synthetically generated or open-source datasets. All the code used to run the experiments is written using open-source libraries or adapted from published code from other papers. We will also release our code and any synthetically generated data to ensure that the work can be reproduced. The human evaluation was approved by the Ethics Committee of the main authors’ employer.
@@ Appendix D Case Studies
      AMR-LDA       The lion is funny
      or the tiger is beautiful.